OpenStudy (anonymous):

Help please! Each base angle of an isosceles triangle is 30 more than twice the measure of the vertex angle. What is the measure of the vertex angle?

OpenStudy (anonymous):

2(30+2x) + x = 180 60+4x+x=180 5x=120 x=24 vertex angle = 24

OpenStudy (anonymous):

thank you :)

OpenStudy (lgbasallote):

Hint: isosceles means the base angles are equal. The sum of the measures of the three angles is 180 degrees so the formula for the angles is: x + y + z = 180 where x, y, and z are angles since it is isosceles.. x= y your base angle is given to be 30 + 2z where z is the vertex angle so x = y = 30 + 2z (30 + 2z) + (30 + 2z) + z = 180

OpenStudy (anonymous):

thanks "]